OpenStudy (anonymous):

copute the integrals ∫ pai cos(2x) dx

OpenStudy (psymon):

\[\int\limits_{}^{}\pi \cos2xdx \] if u = 2x then: du = 2dx dx = du/2 \[\pi \int\limits_{}^{}cosu*\frac{ du }{ 2 } \] \[\frac{ \pi }{ 2 }\int\limits_{}^{}cosudu \implies \frac{ \pi }{ 2 }sinu + C\]